OpenStudy (anonymous):

Find the equation of the line tangent to the graph of the given function at the point with the indicated x-coordinate. f(x) = (x^0.7 + 1)(x^2 + x); x=1

OpenStudy (jhannybean):

There isn't even a point given? just x=1? :\

OpenStudy (jhannybean):

well, first expand f(x) so you can take the derivative..\[f(x)=(x^{0.7}+1)(x^2+x)\]\[f(x) = x^{2.7}+x^{1.7}+x^2+x\]take the derivative..\[f'(x)=2.7x^{1.7}+1.7x^{0.7}+2x+1\] plug in x=1\[f'(x)=2.7(1)^{1.7}+1.7(1)^{0.7}+2(1)+1\]\[f'(x)=7.4\] now you can plug in the slope, 7.4 and x=1 into \[y=mx+b\] to solve for the equation of the tangent line.

OpenStudy (jhannybean):

sorry, your \[f(x)=x^{7.4}\] and then your derivative would be \[f'(x)=7.4 x^{6.4}\]

OpenStudy (jhannybean):

but if you were to plug in x=1... your slope would still = 7.4...

OpenStudy (anonymous):

How do I know what y is when I use the formula to find the tangent line?

OpenStudy (jhannybean):

you have \[y=mx+b\] and you're given x=1 and found your slope = 7.4, but you don't have a point which you could use to plug into y....\[y=7.4(1)+b\] that doesn't make sense to me :\ no point given...

OpenStudy (anonymous):

@jhannybean, the "point" would be \((x,f(x))\). Since \(x=1\), the point you would use is \((1,f(1))=(1,4)\).

OpenStudy (jhannybean):

ohhh... thank you!!! :D i've forgotten. haha.

OpenStudy (jhannybean):

okay,so you've found your m= 7.4and the point (1,4) into y=mx+b 4=(7.4)(1)+b -7.4+4 =b -3.4=b so reform your slope intercept form equation. y=mx+b y=7.4x -3.4

OpenStudy (jhannybean):

Thank you @SithsAndGiggles I was stumped fora minute. haha.

OpenStudy (anonymous):

Ok. Thank you very much!!

OpenStudy (jhannybean):

no problem :)

OpenStudy (anonymous):

@Jhannybean , you're welcome